# MeanCache: User-Centric Semantic Cache for Large Language Model Based Web Services

### v3

@@ 7. Conclusion
MeanCache is the first user-centric semantic cache for Language Model (LLM) based web services, such as ChatGPT. In MeanCache, clients train the global embedding model using federated learning on their local data, while preserving user privacy. The global model formed after the aggregation generates high-quality embeddings for semantic matching. When a new user’s query is duplicated, MeanCache semantically matches the new query with the user’s local cache and returns the most relevant results. This process reduces the computational cost of the LLM service, improves bandwidth and latency, and conserves the user’s query quota for the LLM service. Despite compressing embeddings and saving 83% of storage space, MeanCache outperforms the existing baseline semantic cache (GPTCache). With its distribution cache design, MeanCache offers pathways to reduce one-third of the LLM query inference costs related to semantically similar queries.

### v4

@@ VI Conclusion
MeanCache introduces the first user-centric semantic cache designed for LLM-based web services, such as ChatGPT. In MeanCache, clients collaboratively train a global embedding model using FL on their local data, ensuring user privacy. After aggregation, the global model produces high-quality embeddings for effective semantic matching. When a new query from the user matches a previous one, MeanCache semantically compares it with the user’s local cache and retrieves the most relevant results. This approach reduces the computational cost of LLM services, enhances bandwidth and latency, and conserves the user’s query quota. Even with compressed embeddings that save 83% of storage space, MeanCache outperforms existing baseline. With its distributed cache design, MeanCache offers a solution to reduce up to one-third of LLM query inference costs for semantically similar queries on the user side.
Acknowledgment. We thank anonymous reviewers for providing valuable and constructive feedback to help improve the quality of this work. This work was supported by Amazon - Virginia Tech Initiative in Efficient and Robust Machine Learning. We also thank the Advanced Research Computing Center at Virginia Tech and the Flower FL framework for their support in building and evaluating this work.